### 19 Apr 2026 v2

# Current conservation and amplitude regularisation of the Landau problem: Bohm–Madelung description

Abstract
This work investigates the dynamics of a charged particle in a uniform magnetic field within the Bohm–Madelung formulation of quantum mechanics. In this representation, the stationary Schrödinger equation separates into coupled amplitude and phase equations, where the amplitude sector admits a Sturm–Liouville structure supporting Ermakov–Lewis invariants.
The analysis considers two complementary regularisation schemes: a global Fisher–information–based regularisation and a local canonical (shell) Bohm regularisation derived from stationary flux closure. These are applied within distinct classes of stationary flow, characterised by vanishing and nonvanishing current components.
It is shown that the radial and axial sectors remain globally regularisable, preserving analytic structure across the domain. In contrast, the azimuthal sector develops a nonseparable, generally complex-valued amplitude structure due to gauge-induced coupling. Nevertheless, a consistent local regularity is recovered at the level of canonical branches, where amplitude–momentum relations organise the solution in a well-defined manner.
Regularisation thus acts as a structural reorganisation mechanism in amplitude space, preserving the Landau spectral scale while reorganising the flux-sector structure through branch-wise amplitude–momentum relations, thereby establishing a natural framework for the description of stationary Bohmian dynamics in the Landau problem.
1 Introduction
The quantum mechanics of a charged particle in a magnetic field, commonly referred to as the Landau problem, is well understood and exactly solvable in standard quantum mechanics [1]. The resulting energy spectrum exhibits a harmonic oscillator–like structure with degeneracy in the angular quantum number $l$. The same physical system may be described within the de Broglie–Bohm (Bohm–Madelung) formulation, where the phase $S$ satisfies a Hamilton–Jacobi–type equation supplemented by a self-scaled quantum potential, while the imaginary part enforces a stationary flux condition.
In this formulation, the equations are nonlinearly coupled; however, by utilising the stationary current condition together with amplitude normalisation of the probability density as a constraint, one obtains an exactly solvable structure for the amplitude. In particular, the component-wise separable equations reduce to Sturm–Liouville problems of Helmholtz type [2], and the combined structure naturally yields Ermakov–Pinney–type equations [3, 4, 5]
A key structural feature of this formulation is that the stationary current equation, $\nabla\cdot(\mathbf{Pp})=0$, imposes an amplitude–momentum constraint that organises the admissible solutions. This constraint plays a role analogous to finite flux conditions in classical fluid dynamics, where conservation of mass enforces reciprocal relations between density and velocity. In the Bohmian setting, this manifests as a kinematic condition that regulates the behaviour of momentum near amplitude zeros and underlies the emergence of canonical regularised branches.
Although the resulting equations are exactly solvable, the quantum potential in the radial sector can exhibit singular behaviour due to its explicit dependence on the amplitude. This motivates the introduction of regularisation schemes that yield well-behaved solutions.
The problem of regularisation is not new in quantum mechanics. Langer [6] showed, within the WKB quantisation framework, that an inverse-square regularisation term restores well-behaved solutions without altering the spectral structure. In the Bohm–Madelung formulation, similar inverse-square structures arise naturally in the amplitude equations and obstruct direct analytical treatment.
In recent work [7], regularisation schemes were introduced based on Fisher-information variational principles (global) and canonical closure (local). These schemes render the equations analytically tractable while preserving the spectral structure.
In the Bohm–Madelung formulation, the stationary Landau problem is recast in terms of a Hamilton–Jacobi equation coupled to a continuity equation. These equations are intrinsically nonlinear and coupled, leading to multiple structurally distinct regimes depending on current conditions and regularisation choice.
The purpose of the present work is to analyse how this structure manifests in the Landau problem. The analysis is organised around two classifications: (i) stationary current structure and (ii) regularisation scheme. This leads to a structured classification of solution regimes, summarised in Fig. 1. The primary distinction is set by the stationary current branches ($C_{i}=0$ and $\sum_{i}C_{i}=0,\;C_{i}\neq 0$), which define the underlying solution structure. Within each branch, different regularisation schemes (Ermakov or shell) yield corresponding amplitude constructions.
The programme outlined in Fig. 1 is developed in the following sections. Section 2 establishes the Ermakov–Lewis invariant structure. Section 3 analyses nonlinear behaviour in the azimuthal sector. Section 4 develops global and local regularisation and highlights the contrast between sectors. Section 5 constructs admissible branch solutions. Section 6 discusses representative evolutions of the total amplitudes under different stationary current conditions and their corresponding spectral forms.
The magnetic vector potential introduces a structurally nontrivial coupling in the azimuthal sector, leading to nonlinearity, inseparability, and nonlocal behaviour without invoking hidden variables. This nonlocality arises parametrically from the gauge structure and remains consistent with causality, in line with established analyses [8]. Furthermore, stationary current conservation with nonvanishing components induces a nonlinear amplitude structure exhibiting a Schwarzian–Riccati-type branching in momentum space. Within this framework, local variational principles enable the construction of regularised amplitudes and the identification of admissible branches that support well-behaved stationary dynamics.
Figure 1: Regularisation summary of the Bohm–Madelung description of the Landau problem in the $(R,\Theta,Z)$ sector decomposition.
Bohm–Madelung Form $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ Uniform magnetic field Stationary Current Conditions $C_{i}=0$ $\sum_{i}C_{i}=0$ ; $C_{i}\neq 0$ Ermakov invariant structure Canonical (shell) regularisation Ermakov invariant structure Canonical (shell) regularisation Radial & Axial EP / invariant structure Azimuthal: flux-coupled nonlocal sector Landau spectrum under invariant structure Radial &\& Axial closed-form regularised sectors Azimuthal: simple nonlocal real form Shifted $\sqrt{l^{2}+\tfrac{1}{4}}$ spectrum Radial & Axial: finite-domain regularisable Azimuthal: Whittaker closed form, globally nonreal Energy unchanged from invariant branch Radial & Axial Gaussian-damped branches Azimuthal: branch-wise real amplitude Same shifted $\sqrt{l^{2}+\tfrac{1}{4}}$ spectrum Summary and Conclusion
2 Ermakov–Lewis Invariants structure – a backbone for component current description
A brief account of Ermakov–Lewis–type invariants, which arise naturally in the stationary Bohm–Madelung formulation [5] and in the dynamics of a charged particle in a magnetic field [9], is central to the Landau problem. These invariants characterise the structure underlying the associated differential equations. Extensive treatments of both stationary and time-dependent invariant formulations, including their recasting into Riccati-type equations, are available in [10, 11]. In the present work, we analyse different stationary current conditions and construct the Bohmian amplitude using the corresponding Ermakov–Pinney equations.
2.1 Ermakov–Pinney structure and invariants for the separated sectors
Each separated equation of the stationary problem is of Sturm–Liouville type. After a Liouville transformation, each sector can be written in the canonical form
$y_{i}^{\prime\prime}(q_{i})+\Omega_{i}^{2}(q_{i})\,y_{i}(q_{i})=0,$ (1)
where $\Omega_{i}^{2}(q_{i})$ denotes the effective Sturm–Liouville frequency for the corresponding coordinate sector.
For equations of the form (1) with vanishing component-wise current, there exists an auxiliary amplitude $\sigma_{i}(q_{i})$ satisfying the associated Ermakov–Pinney equation
$\sigma_{i}^{\prime\prime}(q_{i})+\Omega_{i}^{2}(q_{i})\sigma_{i}(q_{i})=\frac{k_{i}}{\sigma_{i}^{3}(q_{i})},$ (2)
with $k_{i}$ a constant determined by the Wronskian of two linearly independent solutions of (1). The coupled system admits the Ermakov–Lewis invariant
$I_{i}=\frac{1}{2}\left[(\sigma_{i}y_{i}^{\prime}-\sigma_{i}^{\prime}y_{i})^{2}+k_{i}\left(\frac{y_{i}}{\sigma_{i}}\right)^{2}\right],\qquad i\in\{r,\theta,z\},$ (3)
with $dI_{i}/dq_{i}=0$, independent of the specific form of $\Omega_{i}^{2}(q_{i})$.
2.2 Ermakov–Pinney structure and invariants in cylindrical coordinates
In cylindrical coordinates, the Ermakov–Pinney structure follows directly from the component-wise continuity equation, which fixes the conserved sectorial fluxes and generates the nonlinear $1/y_{i}^{3}$ terms characteristic of the system.
Continuity first integrals.
For a stationary state with separable amplitude $\rho(r,\theta,z)=R^{2}(r)\Theta^{2}(\theta)Z^{2}(z)$, Eq. (A.6) implies conservation of each component of the probability current,
$\frac{1}{r}\partial_{r}(r\rho v_{r})=\frac{1}{r}\partial_{\theta}(\rho v_{\theta})=\partial_{z}(\rho v_{z})=0,$ (4)
where $v_{i}=(\nabla S-e\mathbf{A})_{i}/m$. Substitution of (4) into the stationary energy equation yields nonlinear amplitude equations of Ermakov–Pinney type in each sector.
Canonical Sturm–Liouville and Ermakov–Pinney forms.
For each separated coordinate sector $x_{i}\in\{r,\theta,z\}$, we write the linear amplitude equation as
$y_{i}^{\prime\prime}+\Omega_{i}^{2}(x_{i})\,y_{i}=0,$ (5)
and its associated Ermakov–Pinney partner as
$\sigma_{i}^{\prime\prime}+\Omega_{i}^{2}(x_{i})\,\sigma_{i}=\frac{c_{i}^{2}}{\sigma_{i}^{3}}.$ (6)
The notation $c_{i}$ denotes integration constants arising from (4), while $C_{i}$ denotes nonzero current components.
Separated equations in cylindrical coordinates.
For $\sqrt{\rho(r,\theta,z)}=R(r)\Theta(\theta)Z(z)$ with the quantisation condition $\partial_{\theta}S=l\hbar$}, the radial, angular, and axial equations (for vanishing current components) take the form
$\chi^{\prime\prime}(r)+\left[\kappa_{r}^{2}-\beta^{2}r^{2}\right]\chi(r)$ $=\frac{c^{2}_{r}}{\chi^{3}(r)},$ (7)
$\Theta^{\prime\prime}(\theta)+\left[l^{2}-2\beta lr^{2}\right]\Theta(\theta)$ $=\frac{c^{2}_{\theta}}{\Theta^{3}(\theta)},$ (8)
$Z^{\prime\prime}(z)+k_{z}^{2}Z(z)$ $=\frac{c^{2}_{z}}{Z^{3}(z)},$ (9)
where $\beta=eB/(2\hbar)$ and the azimuthal equation retains parametric dependence on $r$ due to the gauge coupling.
$\Omega_{r}^{2}(r)$ $=\kappa_{r}^{2}-\beta^{2}r^{2},\qquad\kappa^{2}_{r}=\frac{2mE}{\hbar^{2}}-k_{z}^{2},\qquad\beta=\frac{eB}{2\hbar},$ (10)
$\Omega_{\theta}^{2}(\theta)$ $=l^{2}-2\beta lr^{2},$ (11)
$\Omega_{z}^{2}(z)$ $=k_{z}^{2}.$ (12)
Explicit Ermakov solutions in the separated sectors.
For the radial sector, after Liouville reduction,
$u_{r}^{\prime\prime}+\left(\kappa_{r}^{2}-\beta^{2}r^{2}\right)u_{r}=0,$ (13)
with independent solutions [12]
$u_{r,1}(r)$ $=e^{-\beta r^{2}/2}\,_{1}F_{1}\!\left(-n_{r},\tfrac{1}{2};\beta r^{2}\right),\qquad\kappa_{r}^{2}=\beta(4n_{r}+1),$ (14)
$u_{r,2}(r)$ $=re^{-\beta r^{2}/2}\,{{}_{1}F_{1}}\!\left(-n_{r},\tfrac{3}{2};\beta r^{2}\right),\qquad\kappa_{r}^{2}=\beta(4n_{r}+3).$ (15)
The general radial solution is
$\chi_{r}(r)=\sqrt{A_{r}\,u_{r,1}^{2}(r)+B_{r}\,u_{r,2}^{2}(r)+2D_{r}\,u_{r,1}(r)u_{r,2}(r)},$ (16)
with
$W_{r}=u_{r,1}u_{r,2}^{\prime}-u_{r,2}u_{r,1}^{\prime},\qquad A_{r}B_{r}-D_{r}^{2}=\frac{c^{2}_{r}}{W_{r}^{2}}.$ (17)
The azimuthal and axial amplitudes admit analogous constructions,
$\Theta(\theta)=\sqrt{A_{\theta}\cos^{2}(\Omega_{\theta}\theta)+B_{\theta}\sin^{2}(\Omega_{\theta}\theta)+2D_{\theta}\sin(\Omega_{\theta}\theta)\cos(\Omega_{\theta}\theta)},$ (18)
with
$W_{\theta}=\Omega_{\theta},\qquad A_{\theta}B_{\theta}-D_{\theta}^{2}=\frac{c^{2}_{\theta}}{\Omega_{\theta}^{2}},$ (19)
and
$Z(z)=\sqrt{A_{z}\cos^{2}(k_{z}z)+B_{z}\sin^{2}(k_{z}z)+2D_{z}\sin(k_{z}z)\cos(k_{z}z)},$ (20)
with
$W_{z}=k_{z},\qquad A_{z}B_{z}-D_{z}^{2}=\frac{c^{2}_{z}}{k_{z}^{2}}.$ (21)
The total energy of the system is labeled as $E_{\mathrm{EL}}$ and given by
$\boxed{E_{n_{r}l}(k_{z})=\hbar\omega_{c}\,(n_{r}+\tfrac{1}{2})+\frac{\hbar l}{2m}\left(|eB|-eB\right)+\frac{\hbar^{2}k_{z}^{2}}{2m}},\qquad\omega_{c}=\frac{|e|B}{m}.$ (22)
3 General form of amplitudes for stationary states with nonzero sectorial flux currents
The stationary de Broglie–Bohm continuity equation is
$\nabla\cdot\mathbf{J}=0,\qquad\mathbf{J}=\frac{\rho}{m}\,(\nabla S-e\mathbf{A}).$ (23)
Here, the only condition imposed on the currents is $\bf\nabla\cdot J=0$}, with separable divergence components allowing nonzero sectorial contributions. If we use as a separation constant, then component-wise differential equations are written down as
$\frac{1}{rR^{2}}\frac{d}{dr}\left(rR^{2}p_{r}\right)$ $=C_{r},$ (24)
$\frac{1}{r\Theta^{2}}\frac{d}{d\theta}\left[\Theta^{2}\left(p_{\theta}-\frac{eBr}{2}\right)\right]$ $=C_{\theta},$ (25)
$\frac{1}{Z^{2}}\frac{d}{dz}\left(Z^{2}p_{z}\right)$ $=C_{z}.$ (26)
and ${\bf\nabla\cdot J=0}\text{ }(C_{r}+C_{\theta}+C_{z}=0)$.
For clarity, the conventions and component-wise current relations used throughout this section are summarised in Appendix A.
The general form of separated continuity equations take the Liouville form
$\frac{d}{dq_{i}}\big(w_{i}u_{i}^{2}(p_{i}-eA_{i})\big)=C_{i}\,w_{i}u_{i}^{2},$ (27)
$w_{i}$ are Liouville weights and $u_{i}$ are the respective amplitudes after Liouville scaling. Upon integration, we get
$w_{i}u_{i}^{2}(p_{i}-eA_{i})=K_{i}+C_{i}\int^{q_{i}}w_{i}u_{i}^{2}\,dq_{i},$ (28)
with the global constraint
$\sum_{i}C_{i}=0,$ (29)
ensuring overall current conservation.
For cylindrical coordinates,
$w_{r}=r,\qquad w_{\theta}=w_{z}=1.$
The radial and axial components remain parametrically separable, with amplitudes depending only on their respective coordinates. In contrast, the azimuthal sector retains explicit coupling between $r$ and $\theta$ through the gauge term $A_{\theta}$, and is therefore not reducible to a purely single-variable description.
The azimuthal component is therefore of primary interest in this work, as it exhibits nonlinearity and coordinate-wise nonlocality induced by the gauge coupling. The stationary current and energy equations for the azimuthal sector are
$\frac{dp_{\theta}}{d\theta}+\left(p_{\theta}-\frac{eBr}{2}\right)\frac{d}{d\theta}\ln\Theta^{2}=rC_{\theta},$ (30)
$\frac{1}{2m}\left(p_{\theta}^{2}-eBr\,p_{\theta}\right)-\frac{\hbar^{2}}{2mr^{2}}\frac{\Theta^{\prime\prime}}{\Theta}=\frac{\hbar^{2}l^{2}}{2mr^{2}}.$ (31)
Using the Hamilton–Jacobi representation $p_{\theta}=\partial_{\theta}S/r$, the gauge-induced $B$-dependence becomes explicit in the azimuthal dynamics.
A useful reduction is obtained by introducing the logarithmic derivative
$w(\theta):=\frac{\Theta^{\prime}(\theta)}{\Theta(\theta)}.$ (32)
Then
$\frac{\Theta^{\prime\prime}}{\Theta}=w^{\prime}+w^{2}.$ (33)
Using the shifted angular momentum variable
$\pi_{\theta}(\theta):=p_{\theta}-\frac{eBr}{2},$ (34)
the local angular system becomes
$\pi_{\theta}^{\prime}+2w\pi_{\theta}=rC_{\theta},$ (35)
and from the energy term,
$\pi_{\theta}^{2}=\hbar^{2}(w^{\prime}+w^{2})+\Lambda_{\theta},\qquad\Lambda_{\theta}:=l^{2}+\frac{e^{2}B^{2}r^{2}}{4}.$ (36)
Eqs. (35) and  (36) form the coupled first-order nonlinear system  (37).
$\pi_{\theta}^{\prime}$ $=rC_{\theta}-2w\pi_{\theta},$ (37)
$w^{\prime}$ $=\frac{\pi_{\theta}^{2}-\Lambda_{\theta}}{\hbar^{2}}-w^{2}.$
Eliminating $w(\theta)$, one obtains a second-order nonlinear differential equation for $\pi_{\theta}(\theta)$, given by
$3\pi_{\theta}^{\prime 2}-2\pi_{\theta}\pi_{\theta}^{\prime\prime}-4r^{2}C_{\theta}\pi_{\theta}^{\prime}-\frac{4}{\hbar^{2}}\pi_{\theta}^{4}+4\Lambda_{\theta}\pi_{\theta}^{2}+r^{4}C_{\theta}^{2}=0.$ (38)
If the first-order system (37) is solved, the logarithmic derivative $w(\theta)$ can be obtained and the angular amplitude reconstructed via
$\Theta(\theta)=\Theta_{0}\exp\!\left(\int^{\theta}w(\varphi)\,d\varphi\right).$ (39)
Thus, the amplitude may be recovered either from the $(\pi_{\theta},w)$ system or, equivalently, from the reduced $\pi_{\theta}(\theta)$ equation. In this representation, the second-order ratio $\Theta^{\prime\prime}/\Theta$ is replaced by a Riccati-type structure in $w(\theta)$, which is convenient for analysing the nonlinear behaviour of the azimuthal sector.
3.1 Amplitude function for $\partial_{\theta}S\neq l\hbar$ and $C_{\theta}=0$
The assumption $\partial_{\theta}S=l\hbar$ is a restrictive sector of azimuthal current condition for zero current in the $\hat{\theta}$ direction that is well understood in quantum mechanics. In the Bohmian description, we remove this condition to obtain phase space dynamics governed by continuity relations. The choice reveals a rich class of solutions and global symmetries through first integrals. Using
$\frac{\partial_{\theta}S}{r}=p_{\theta}$
and the constant current condition
$\frac{d}{d\theta}\left[\Theta^{2}\left(\partial_{\theta}S-\hbar\frac{\Phi}{\Phi_{0}}\right)\right]=0$
${\partial_{\theta}S}=\left(\frac{\kappa_{\theta}}{\Theta^{2}}+\hbar\frac{\Phi}{\Phi_{0}}\right),\quad\kappa_{\theta}\text{ is an integration constant.}$
The azimuthal equation
$\frac{p_{\theta}^{2}}{2m}-\frac{eBr}{2m}p_{\theta}-\frac{\hbar^{2}}{2mr^{2}}\frac{\Theta^{\prime\prime}}{\Theta}=\frac{\hbar^{2}l^{2}}{2mr^{2}}$
becomes
$\left(\frac{\kappa_{\theta}}{\Theta^{2}}+\hbar\frac{\Phi}{\Phi_{0}}\right)^{2}-\hbar\frac{\Phi}{\Phi_{0}}\left(\frac{\kappa_{\theta}}{\Theta^{2}}+\hbar\frac{\Phi}{\Phi_{0}}\right)-\hbar^{2}\frac{\Theta^{\prime\prime}}{\Theta}$ $=\hbar^{2}l^{2}$ (40)
$\Rightarrow\quad\frac{\kappa_{\theta}^{2}}{\Theta^{4}}+\hbar\frac{\Phi}{\Phi_{0}}\frac{\kappa_{\theta}}{\Theta^{2}}-\hbar^{2}\frac{\Theta^{\prime\prime}}{\Theta}$ $=\hbar^{2}l^{2}$ (41)
$\Theta^{\prime\prime}+l^{2}\Theta-\frac{\kappa_{\theta}}{\hbar}\frac{\Phi}{\Phi_{0}}\frac{1}{\Theta}-\frac{\kappa_{\theta}^{2}}{\hbar^{2}}\frac{1}{\Theta^{3}}=0.$ (42)
Upon multiplying Eq. (42) by $\Theta^{\prime}$ and integrating once with respect to $\theta$ we obtain the first integral
$\frac{1}{2}(\Theta^{\prime})^{2}+\frac{1}{2}l^{2}\Theta^{2}-\frac{\kappa_{\theta}}{\hbar}\frac{\Phi}{\Phi_{0}}\ln|\Theta|+\frac{\kappa_{\theta}^{2}}{2\hbar^{2}\Theta^{2}}=\mathcal{E}_{\theta},$ (43)
where $\mathcal{E}_{\theta}$ is the integration constant.
Equivalently,
$(\Theta^{\prime})^{2}=2\mathcal{E}_{\theta}-l^{2}\Theta^{2}+2\frac{\kappa_{\theta}}{\hbar}\frac{\Phi}{\Phi_{0}}\ln|\Theta|-\frac{\kappa_{\theta}^{2}}{\hbar^{2}\Theta^{2}}.$ (44)
Hence the implicit solution is given by the quadrature
$\boxed{\theta-\theta_{0}=\pm\int\frac{d\Theta}{\sqrt{2\mathcal{E}_{\theta}-l^{2}\Theta^{2}+2\frac{\kappa_{\theta}}{\hbar}\frac{\Phi}{\Phi_{0}}\ln|\Theta|-\frac{\kappa_{\theta}^{2}}{\hbar^{2}\Theta^{2}}}}}$ (45)
From the above expressions, we notice a first integral similar to the Ermakov-Lewis invariant but it contains a $ln|\Theta|$ term coupled to flux as a perturbation or symmetry breaking term, further first integral of $\Theta$ does not have a closed form representation in its coordinate description. In  3.2 we show an alternate route to obtain a closed form of $\Theta$ solution by changing the basis of representation from amplitude-coordinate basis to an amplitude-momentum basis.
3.2 $\Theta(\theta)$ through $\pi_{\theta}(\theta)$, change of basis, closed form expression for $C_{\theta}=0$
For the zero-current branch $C_{\theta}=0$, the nonlinear equation for the shifted angular momentum
$\pi_{\theta}(\theta):=\partial_{\theta}S-\hbar\frac{\Phi}{\Phi_{0}}$ (46)
reduces Eq. (31) to
$3\pi_{\theta}^{\prime 2}-2\pi_{\theta}\pi_{\theta}^{\prime\prime}-\frac{4}{\hbar^{2}}\pi_{\theta}^{4}+4\Lambda_{\theta}\pi_{\theta}^{2}=0,\qquad\Lambda_{\theta}=l^{2}+\left(\frac{\Phi}{\Phi_{0}}\right)^{2}.$ (47)
Introducing
[…]
$\xi^{2}-2\pi_{\theta}\,\xi\,\frac{d\xi}{d\pi_{\theta}}-\frac{4}{\hbar^{2}}\pi_{\theta}^{2}+4\Lambda_{\theta}=0.$
With another substitution for $\xi^{2}(\pi_{\theta})=Y$, we get
$Y-\pi_{\theta}\frac{dY}{d\pi_{\theta}}-\frac{4}{\hbar^{2}}{\pi^{2}_{\theta}}+4\Lambda_{\theta}=0$ (48)
This is a linear differential equation of $Y(\pi_{\theta})$, which upon using an integrating factor $1/\pi_{\theta}$
$\frac{\pi^{\prime}_{\theta}}{\pi_{\theta}}=\pm\sqrt{\mathcal{E}_{\pi}\pi_{\theta}-\frac{4}{\hbar^{2}}\pi_{\theta}^{2}-4\Lambda_{\theta}}$ (49)
where $\mathcal{E_{\pi}}$ is an integration constant. Therefore the implicit solution is
$\theta-\theta_{0}=\pm\int\frac{d\pi_{\theta}}{\pi_{\theta}\sqrt{\mathcal{E}_{\pi}\pi_{\theta}-\frac{4}{\hbar^{2}}\pi_{\theta}^{2}-4\Lambda_{\theta}}}.$ (50)
Since the denominator is a quadratic polynomial in $\pi_{\theta}$, this integral is reducible to elementary trigonometric or hyperbolic functions, depending on the sign of the discriminant.
Define the discriminant
$\Delta_{\pi}:=\mathcal{E}_{\pi}^{2}-\frac{64\Lambda_{\theta}}{\hbar^{2}}.$ (51)
The $\Lambda_{\theta}$ is a nonnegative number and one obtains for $\Lambda_{\theta}>0$
$\theta-\theta_{0}=\pm\frac{1}{2\sqrt{\Lambda_{\theta}}}\sin^{-1}\!\left(\frac{\frac{8\Lambda_{\theta}}{\pi_{\theta}}-\mathcal{E}_{\pi}}{\sqrt{\Delta_{\pi}}}\right),\qquad\Delta_{\pi}>0,$ (52)
equivalently
$\boxed{\pi_{\theta}(\theta)=\frac{8\Lambda_{\theta}}{\mathcal{E}_{\pi}+\sqrt{\Delta_{\pi}}\,\sin\!\bigl(2\sqrt{\Lambda_{\theta}}(\theta-\theta_{0})\bigr)}.}$ (53)
3.3 Azimuthal nonlocal structure – a consequence of magnetic vector potential
Let us analyse a simple case for zero current branch. For the branch $C_{\theta}=0$, the shifted angular momentum is
$\pi_{\theta}(\theta):=\partial_{\theta}S-\hbar\frac{\Phi}{\Phi_{0}}=\frac{8\Lambda_{\theta}}{\mathcal{E}_{\pi}+\sqrt{\Delta_{\pi}}\,\sin\!\bigl(2\sqrt{\Lambda_{\theta}}(\theta-\theta_{0})\bigr)},$ (54)
where
$\Delta_{\pi}=\mathcal{E}_{\pi}^{2}-\frac{64\Lambda_{\theta}}{\hbar^{2}},\qquad\Lambda_{\theta}=l^{2}+\left(\frac{\Phi}{\Phi_{0}}\right)^{2}.$ (55)
Hence
$\partial_{\theta}S=\hbar\frac{\Phi}{\Phi_{0}}+\frac{8\Lambda_{\theta}}{\mathcal{E}_{\pi}+\sqrt{\Delta_{\pi}}\,\sin\!\bigl(2\sqrt{\Lambda_{\theta}}(\theta-\theta_{0})\bigr)}.$ (56)
Integrating with respect to $\theta$, one finds
$\boxed{S_{\theta}(\theta)=\hbar\frac{\Phi}{\Phi_{0}}\,\theta+\hbar\,\tan^{-1}\!\left(\frac{\mathcal{E}_{\pi}\,\tan\!\bigl(\sqrt{\Lambda_{\theta}}(\theta-\theta_{0})\bigr)+\sqrt{\Delta_{\pi}}}{8\sqrt{\Lambda_{\theta}}/\hbar}\right)+S_{\theta,0},}$ (57)
where $S_{\theta,0}$ is an integration constant.
The azimuthal sector in the Bohm–Madelung formulation of the Landau problem is not genuinely separable once magnetic flux is retained. As seen from above, special case $C_{\theta}=0$, the equation admits a first integral and becomes separable in its form, but retains radial dependence through $\Lambda_{\theta}$, the nonlinear coupling, and the integration constant. This demonstrates that the nonlocality present is structural, arising from the coupling of kinematic variables through the vector potential, rather than from hidden-variable interactions. The resulting framework differs from standard Bohmian ontology, as the inseparability originates at the level of the differential structure of the equations, while preserving causal determinism.
3.4 Nonlinearity structure of azimuthal equation for $C_{\theta}\neq 0$
For representation purposes, it is useful to reduce the Eq. (38) to a nonlinear first order differential equation through new variables. Through a sequence of change of variables, it is possible to illustrate the intrinsic nonlinearity arising from the mixed term of $\theta$ and $r$ due to magnetic vector potential.
Recalling the vector potential and logarithmic derivative of the amplitude $\Theta$
$\pi_{\theta}(\theta):=\partial_{\theta}S-\hbar\frac{\Phi}{\Phi_{0}},\qquad w(\theta):=\frac{\Theta^{\prime}}{\Theta}.$
We begin with Eq. (38)
$3\pi_{\theta}^{\prime 2}-2\pi_{\theta}\pi_{\theta}^{\prime\prime}-4r^{2}C_{\theta}\pi_{\theta}^{\prime}-\frac{4}{\hbar^{2}}\pi_{\theta}^{4}+4\Lambda_{\theta}\pi_{\theta}^{2}+r^{4}C_{\theta}^{2}=0.$
and introduce another variable $\xi(\pi_{\theta})$,
$\xi(\pi_{\theta}):=\frac{\pi_{\theta}^{\prime}}{\pi_{\theta}}.$
Then, with a bit of algebra and re-arrangement we get
$\xi^{2}-2\pi_{\theta}\,\xi\,\frac{d\xi}{d\pi_{\theta}}-4r^{2}C_{\theta}\frac{\xi}{\pi_{\theta}}-\frac{4}{\hbar^{2}}\pi_{\theta}^{2}+4\Lambda_{\theta}+\frac{r^{4}C_{\theta}^{2}}{\pi_{\theta}^{2}}=0.$ (58)
A further reduction is obtained by defining
$F(\pi_{\theta}):=\xi^{2}\pi_{\theta}.$
Since
$\frac{dF}{d\pi_{\theta}}=\xi^{2}+2\pi_{\theta}\,\xi\,\frac{d\xi}{d\pi_{\theta}},$
it follows that
$\xi^{2}-2\pi_{\theta}\,\xi\,\frac{d\xi}{d\pi_{\theta}}=\frac{2F}{\pi_{\theta}}-\frac{dF}{d\pi_{\theta}}.$
Substituting this into (58) gives
$\boxed{\frac{dF}{d\pi_{\theta}}=\frac{2F}{\pi_{\theta}}\mp\frac{4r^{2}C_{\theta}}{\pi_{\theta}}\sqrt{\frac{F}{\pi_{\theta}}}-\frac{4}{\hbar^{2}}\pi_{\theta}^{2}+4\Lambda_{\theta}+\frac{r^{4}C_{\theta}^{2}}{\pi_{\theta}^{2}},}$ (59)
where the sign corresponds to the two branches
$\xi=\pm\sqrt{\frac{F}{\pi_{\theta}}}.$ (60)
Equation (59) is the desired first-order nonlinear equation in $\pi_{\theta}$-space. In the zero-current limit $C_{\theta}=0$, the square-root term disappears and the equation reduces to the linear flow
$\frac{dF}{d\pi_{\theta}}-\frac{2}{\pi_{\theta}}F=-\frac{4}{\hbar^{2}}\pi_{\theta}^{2}+4\Lambda_{\theta},$ (61)
which reproduces the reducible Ermakov branch whose solution is Eq. (50). By contrast, for $C_{\theta}\neq 0$, the term
$\mp\frac{4r^{2}C_{\theta}}{\pi_{\theta}}\sqrt{\frac{F}{\pi_{\theta}}}$ (62)
introduces a genuine square-root branching, while the term
$\frac{r^{4}C_{\theta}^{2}}{\pi_{\theta}^{2}}$ (63)
provides the accompanying nonlinear current correction thus making it irreducible. Thus the nonzero-current azimuthal branch is naturally interpreted as a square-root deformation of the zero-current linear flow. The branching originates from the nonlinear $r^{2}C_{\theta}$- dependence of the principal branch
$\xi=\pm\sqrt{\frac{F}{\pi_{\theta}}}=\pm\frac{\pi^{\prime}_{\theta}}{\pi_{\theta}},$ (64)
so that the loss of Ermakov invariance is directly tied to the current-induced branching structure in momentum space. The passage to the variable $F$ is introduced only to exhibit this nonlinearity in the simplest first-order form. It may further be shown that the second-order equation for $\pi_{\theta}$, as well as the corresponding radial and axial equations for nonzero $C_{i}$, can be recast as projective Schwarzian differential equations for the associated canonical Hamilton–Jacobi actions $S_{i}$. In that representation, however, all three component equations take the same formal form, and the distinctive azimuthal nonlinearity and nonlocality are no longer manifest. The present formulation therefore provides a complete closure directly in terms of the physical variables, without the need to introduce an auxiliary mathematical object solely to obtain the Schwarzian form.
Remark on projective (Schwarzian) structure.
While the nonzero current branches admit a reformulation in terms of Schwarzian differential equations for the canonical Hamilton–Jacobi actions, a complete projective analysis is beyond the scope of the present work.
4 Local regularisation, global extension, and branch structure of the azimuthal sector
In this section, we analyse the stationary system with nonvanishing current components, where the amplitude and momentum equations are nonlinearly coupled at the branch level. This coupling introduces near-singular behaviour that renders the system analytically intractable in its native form. To address this, we first introduce a local regularisation scheme that restores a controlled amplitude–momentum structure and yields analytically tractable solutions in a restricted domain.
A natural question is whether these locally regularised solutions admit a consistent global extension. As shown below, the radial and axial sectors extend globally with modified spectral structure, while the azimuthal sector develops a nonreal-valued amplitude due to gauge-induced coupling. This obstruction necessitates a branch-wise regularisation of the azimuthal amplitude, which is treated separately in Section 5.
Working in cylindrical coordinates with the Landau gauge, we impose the canonical Bohmian regularisation  [7] condition
$q_{i}p_{i}=\frac{\hbar}{2},$ (65)
on the Hamilton–Jacobi variables. With the separable amplitude
$\rho(r,\theta,z)=R^{2}(r)\,\Theta^{2}(\theta)\,Z^{2}(z),$ (66)
the stationary Bohm–Madelung equations reduce, under the combined action of continuity and regularisation constraints, to component-wise ordinary differential equations containing inverse-square terms. The resulting energy constraint takes the form
$\frac{\hbar^{2}}{8mr^{2}}+Q_{r}+\frac{\hbar^{2}}{2m}\left(\frac{1}{2r\theta}-\frac{eBr}{2\hbar}\right)^{2}+Q_{\theta}+\frac{\hbar^{2}}{8mz^{2}}+Q_{z}=E.$ (67)
Here $Q_{i}$ denote the component quantum potentials as defined in (A.11). The energy equation can then be separated and, together with the corresponding component current relations, yields modified Ermakov–Pinney-type equations in each sector.
4.1 Global extension of regularised radial and axial sectors
The radial component of the energy equation becomes a shifted equation of a central field problem
$R^{\prime\prime}+\frac{1}{r}R^{\prime}+\left[\kappa^{2}_{r}-\beta^{2}r^{2}-\frac{(l^{2}+\tfrac{1}{4})}{r^{2}}\right]R=0,$ (68)
with regular solutions  [12]. After a Liouville scaling to reduce the differential for $R(r)$ to a normal form using the transformation $R(r)=\chi_{r}/\sqrt{r}$ and simplification, one obtains a Langer-corrected equation of $\nu$
$\chi_{r}^{\prime\prime}+\left[\kappa_{r}^{2}-\beta^{2}r^{2}-\frac{\nu^{2}-\tfrac{1}{4}}{r^{2}}\right]\chi_{r}=0,\qquad\qquad\nu=\sqrt{l^{2}+\tfrac{1}{4}}$ (69)
Thus the regular solution is
$\chi_{n_{r}\nu}(r)=Cr^{|\nu|+\frac{1}{2}}e^{-\beta r^{2}/2}{{}_{1}}F_{1}(a_{r},\nu+1,\beta r^{2}).$ (70)
Or,
$R_{n_{r}\nu}(r)=Cr^{|\nu|}e^{-\beta r^{2}/2}{{}_{1}}F_{1}(a_{r},\nu+1,\beta r^{2}).$ (71)
where the quantisation condition is satisfied by
$\nu+\frac{1}{2}-\frac{\kappa_{r}^{2}}{4\beta}=a_{r}=-n_{r},\qquad n_{r}\text{ is a nonnegative integer.}$
The analytical form of radial solution is well defined for all values of $r$ and thus we say that, the regularity is maintained throughout the range of $r$.
4.1.1 Axial sector: global extension
The axial equation reads
$Z^{\prime\prime}+\frac{1}{4z^{2}}Z=k_{z}^{2}Z,$ (72)
with regular solution
$Z(z)=\sqrt{z}\,J_{1/\sqrt{2}}(k_{z}z).$ (73)
This equation is identical in structure to the regularised free-particle problem, with the inverse-square term fixing the admissible local behaviour and analytically extendable to all values of $|z|$.
The radial and axial sectors thus admit globally well-defined analytic extensions under canonical regularisation. It is therefore natural to examine whether the azimuthal sector admits a similar global continuation. As shown below, the presence of gauge-induced coupling leads to a fundamentally different behaviour.
The next section examines the azimuthal amplitude under canonical regularisation.
4.1.2 Azimuthal sector: obstruction to global extension
The angular equation obtained after canonical regularisation takes the form
$\Theta^{\prime\prime}+\left(l^{2}+\frac{\varphi}{\theta}-\frac{1}{4\theta^{2}}\right)\Theta=0,$ (74)
where $\varphi=eBr^{2}/(2\hbar)$. This equation may be reduced to the canonical Whittaker differential equation [13]
$\Theta_{xx}+\left(-\frac{1}{4}+\frac{\kappa}{x}+\frac{\frac{1}{4}-\mu^{2}}{x^{2}}\right)\Theta=0,\qquad\mu=\frac{1}{\sqrt{2}}\quad\text{ and }x=2il\theta.$ (75)
Hence the general solution for the angular amplitude is
$\Theta(\theta)=C_{1}\,M_{\kappa,\mu}(2il\theta)+C_{2}\,W_{\kappa,\mu}(2il\theta),\qquad\kappa=-\frac{i}{2l}\frac{\Phi}{\Phi_{0}},\quad\Phi=\pi Br^{2},\quad\Phi_{0}=\frac{h}{e}.$ (76)
Thus the azimuthal amplitude $\Theta(\theta)$ is generically complex. Consequently, the assumption of a globally defined real-valued amplitude leads to a reductio ad absurdum in the presence of gauge-induced coupling.
This establishes that, unlike the radial and axial sectors, the azimuthal component does not admit a globally real separable amplitude and requires a distinct branch-wise regularisation, which is developed in Section 5.
5 Branch regularisation of the azimuthal amplitude
The reductio ad absurdum of a globally real azimuthal amplitude necessitates the preservation of a branch-wise regularisation. The mixed term proportional to $r^{2}/\theta$, arising from the gauge coupling, is the source of the resulting complex branching structure. In the limit $r\to 0$, this mixed term vanishes, and the azimuthal energy equation retains regularity for all values of $|\theta|$.
A complementary local description may be obtained by analysing the general current equation in the regularisation limit $\theta\to 0$, where the branch structure becomes well-defined. We illustrate this by constructing solutions of the current equation for the $\Theta$-branch.
For $\Theta$, with nonvanishing $C_{\theta}$ and the canonical relation $q_{i}p_{i}=\hbar/2$, we obtain
$\frac{\partial S}{\partial\theta}=\frac{\hbar}{2\theta},\qquad p_{\theta}=\frac{1}{r}\frac{\partial S}{\partial\theta}=\frac{\hbar}{2r\theta},$ (77)
$\frac{d}{d\theta}\left[\Theta^{2}\left(p_{\theta}-\frac{eBr}{2}\right)\right]=rC_{\theta}\,\Theta^{2}.$ (78)
Writing the equation in flux-scaled form, we obtain
$\frac{d}{d\theta}\left[\Theta^{2}\left(\frac{1}{2\theta}-\frac{\Phi}{\Phi_{0}}\right)\right]=\frac{r^{2}C_{\theta}}{\hbar}\Theta^{2}.$ (79)
Defining
$\varphi=\frac{\Phi}{\Phi_{0}},\qquad\kappa=\frac{r^{2}C_{\theta}}{\hbar},$
the equation becomes
$\frac{d}{d\theta}\ln\Theta^{2}=\frac{\frac{1}{\theta}+2\kappa\theta}{1-2\varphi\theta}.$ (80)
Integrating,
$\ln\Theta^{2}=\ln|\theta|-\left(1+\frac{\kappa}{2\varphi^{2}}\right)\ln|1-2\varphi\theta|-\frac{\kappa}{\varphi}\theta+\mathrm{const},$ (81)
so that
$\Theta^{2}(\theta)=A_{\theta}\,|\theta|\,|1-2\varphi\theta|^{-\left(1+\kappa/(2\varphi^{2})\right)}\exp\!\left(-\frac{\kappa}{\varphi}\theta\right).$ (82)
Equivalently,
$\boxed{\Theta(\theta)=\sqrt{A_{\theta}}\,|\theta|^{1/2}|1-2\varphi\theta|^{-\frac{1}{2}\left(1+\kappa/(2\varphi^{2})\right)}\exp\!\left(-\frac{\kappa}{2\varphi}\theta\right).}$ (83)
The expression is constructed to be real-valued by selecting an appropriate branch. It exhibits singular behaviour at $\theta=0$ and at $1-2\varphi\theta=0$. The first corresponds to the canonical inverse-square regularisation point, while the second defines a flux-controlled singularity where the azimuthal momentum $p_{\theta}$ diverges. This latter point is governed by the flux ratio $\varphi=\Phi/\Phi_{0}$, and through $\Phi=\pi Br^{2}$, it encodes an implicit dependence on the magnetic length scale.
The solution is restricted to real branches, excluding complex continuations through these singularities, which are not physically admissible under stationary current conditions. The exponential term contains the constant $\kappa$, related to the azimuthal current parameter $C_{\theta}$. Imposing $C_{\theta}\geq 0$ and the stationary constraint $\sum_{i}C_{i}=0$ requires the radial and axial current constants to satisfy $C_{r}+C_{z}\leq 0$. Thus, the admissible angular amplitude is selected by the current branch, and the regularisation in $\theta$ emerges as a consequence of flux-balanced stationary flow rather than an imposed boundary condition.
6 Physical evolution of branch-regularised amplitudes
Building on the branch-wise regularisation developed in Section 5, we examine the admissible physical evolution of the amplitudes $R(r)$, $\Theta(\theta)$, and $Z(z)$ under stationary current constraints. For the nonvanishing radial current sector, prior to regularisation, the radial amplitude satisfies
$R^{\prime\prime}+\frac{1}{r}R^{\prime}+\left[k_{0}^{2}-\beta^{2}r^{2}-\frac{l^{2}}{r^{2}}\right]R=0,$
together with
$\frac{1}{rR^{2}}\frac{d}{dr}(rR^{2}p_{r})=C_{r}.$
Under canonical regularisation one imposes the local closure
$p_{r}\sim\frac{\hbar}{2r},$
so that the continuity equation no longer determines $p_{r}$ freely, but instead constrains the amplitude according to
$R^{2}(r)\propto\exp\!\left(\frac{C_{r}}{\hbar}r^{2}\right).$
Thus the regularised branch defines a constrained sector of the stationary solution space rather than a generic nonzero-current configuration.
For $C_{r}<0$, this yields a finite Gaussian-decaying radial density, compatible with cylindrical normalisability. In the separated stationary-flow picture, the branch constants may be chosen subject to $C_{r}+C_{\theta}+C_{z}=0$, so that radial damping is balanced by the remaining sectors.
Next, for $C_{z}<0$, the regularised amplitude of $Z$ can be shown through the continuity equation
$p_{z}(z)\sim\frac{\hbar}{2z},\qquad p_{z}^{\prime}+2\frac{Z^{\prime}}{Z}p_{z}=C_{z},$ (84)
$\frac{Z^{\prime}}{Z}=\frac{1}{2z}+\frac{C_{z}}{\hbar}z,$ (85)
Thus, in the canonical regularising limit $p_{z}\sim\hbar/(2z)$, the nonvanishing axial branch current $C_{z}$ does not alter the universal square-root behaviour of the amplitude at the node, but deforms it by a Gaussian factor; for $C_{z}<0$ this produces a locally damped branch profile, $Z(z)\sim|z|^{1/2}\exp(-|C_{z}|z^{2}/2\hbar).$
The condition
$2\frac{\Phi}{\Phi_{0}}\theta=1$
defines a flux-dependent angular scale at which the regularised local angular branch $\partial_{\theta}S=\hbar/(2\theta)$ balances the magnetic shift. The resulting singularity is therefore not a fundamental geometric divergence, but a dynamical feature of the reduced continuity equation, reflecting the competition between local canonical regularisation and the flux-induced angular drift. A natural regularising branch choice is to take
$C_{r}<0,\qquad C_{z}<0,\qquad C_{\theta}=-(C_{r}+C_{z}),$ (86)
so that
$C_{r}+C_{\theta}+C_{z}=0.$ (87)
With this assignment, the radial and axial sectors both acquire locally damping branch profiles, while the azimuthal sector carries the compensating current required by the stationary continuity constraint. In particular, the axial canonical branch
$p_{z}\sim\frac{\hbar}{2z}$ (88)
gives
$Z(z)\sim|z|^{1/2}\exp\!\left(-\frac{|C_{z}|}{2\hbar}z^{2}\right)\qquad(C_{z}<0),$ (89)
and an analogous damping behaviour holds in the radial sector for $C_{r}<0$. Thus the separated amplitude may be rendered globally well behaved across the $r$- and $z$-branches.
Finally, the azimuthal expression at the regularisation limit takes the form,
$\Theta(\theta)=\sqrt{A_{\theta}}\,|\theta|^{1/2}|1-2\varphi\theta|^{-\frac{1}{2}\left(1+\kappa/(2\varphi^{2})\right)}\exp\!\left(-\frac{\kappa}{2\varphi}\theta\right).$ (90)
In the azimuthal sector, a flux-induced singularity persists at
$\frac{2\Phi}{\Phi_{0}}\,\theta=1,$ (91)
The azimuthal branch develops a singularity at $(2\Phi/\Phi_{0})\theta=1$, which can take continuous values as functions of $B$ and $r$. In this sense, the nontrivial singular structure of the nonzero-current branch is not generated by the canonical regularisation itself, but arises from the nonlocal coupling induced by the mixed gauge term. Finally, $\kappa$ of the exponent terms of $\Theta(\theta)$ carries the same sign of $C_{\theta}$ and for finiteness of the amplitude, it forces $C_{\theta}\geq 0$.
Further, the continuity condition
$C_{r}+C_{\theta}+C_{z}=0$ (92)
is necessary but not sufficient for full local branch regularity. If regularity is imposed componentwise, one is naturally led to the choice
$C_{r}<0,\qquad C_{z}<0,\qquad C_{\theta}=-(C_{r}+C_{z})>0,$ (93)
so that the radial and axial sectors are individually damping. However, if one instead requires only that the full separated amplitude remain finite, then a local squeezing between the $r$- and $z$-branches may provide the necessary compensation. In that case the angular sector must remain undeformed, and the only consistent choice is
$C_{\theta}=0,\qquad C_{z}=-C_{r}.$ (94)
Thus full local branch regularity may be realised either by componentwise damping or, more economically, by a radial-axial compensating branch with vanishing azimuthal current.
6.1 Energy spectrum at regularisation limit
As a consequence of branch regularisation, the energy spectrum is determined primarily by the radial sector, which remains analytically well-defined under global extension. The axial contribution retains its free-particle character, while the azimuthal sector, despite its nonlocal and branch-dependent amplitude structure, does not alter the spectral form. The total energy of the shell regularised Landau problem is identical in its functional form but with a subleading correction through the Langer-type shift $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$, which becomes negligible for large $l$.
$E_{n_{r}\nu}(k_{z})=E_{\perp}(n_{r},\nu)+E_{z}=\frac{\hbar\omega_{c}}{2}\,(2n_{r}+\nu+1)+\frac{\hbar^{2}k_{z}^{2}}{2m},\qquad\omega_{c}=\frac{|e|B}{m}.$ (95)
Upon substituting $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$, one obtains the canonically regularised energy $E_{\mathrm{CBR}}$ as
$\boxed{E_{n_{r}l}(k_{z})=E_{\perp}(n_{r},l)+E_{z}=\frac{\hbar\omega_{c}}{2}\left(2n_{r}+\sqrt{l^{2}+\tfrac{1}{4}}+1\right)+\frac{\hbar^{2}k_{z}^{2}}{2m}}\qquad\omega_{c}=\frac{|e|B}{m}.$ (96)
In particular, the local regularisation or loss of global analyticity in the azimuthal amplitude does not affect the canonical Bohm-regularised spectrum $E_{\mathrm{CBR}}$. The only modification arises from the radial sector through the Langer-type shift $\nu=\sqrt{l^{2}+\tfrac{1}{4}}$.
Remark on spectral ordering.
For identical radial and axial quantum numbers, one finds $E_{\mathrm{QM}}\leq E_{\mathrm{EL}}\leq E_{\mathrm{CBR}}$ for $l\geq 1$. where, $_{\mathrm{QM}}, EE$$_{\mathrm{EL}}$ and $E_{\mathrm{CBR}}$ are energy spectra obtained from standard quantum mechanics, Ermakov-Lewis and canonical Bohm regularisation routes respectively.
The regularised spectrum separates naturally into a cyclotron contribution and an angular term,
$E_{n_{r}l}=\hbar\omega_{c}\left(n_{r}+\tfrac{1}{2}\right)+\frac{\hbar\omega_{c}}{2}\sqrt{l^{2}+\tfrac{1}{4}}+\cdots$ (97)
The second term lifts the degeneracy with respect to the azimuthal index $l$, which is otherwise present in the standard Landau problem. This contribution arises from the regularised angular structure and may be viewed as a non-cyclotronic contribution induced by canonical regularisation.
7 Conclusion
A consistent canonical Bohm regularisation of a charged particle in a uniform magnetic field leads to a clear structural decomposition of the stationary solution space. The radial and axial sectors admit globally well-defined analytic extensions, while the azimuthal sector, due to gauge-induced coupling, does not admit a globally real separable amplitude. This obstruction is resolved through a branch-wise regularisation, which provides a consistent local description of the azimuthal dynamics.
The analysis shows that the nonlinear coupling introduced by nonvanishing current components leads naturally to a branching structure in amplitude–momentum space. In this framework, regularisation is not imposed as an external condition but emerges as a consequence of stationary current conservation and canonical closure. The resulting structure preserves the Ermakov–Lewis invariant framework while extending it to nontrivial gauge-coupled systems.
Despite the nonlocal and branch-dependent nature of the azimuthal amplitude, the energy spectrum remains governed by the radial sector and retains the canonical Bohm-regularised form $E_{\mathrm{CBR}}$, with only a Langer-type correction. Thus, the spectral structure remains consistent with standard quantum mechanics in the appropriate limit.
Recent experimental observations of free-electron Landau states have revealed structured current flow patterns that are naturally interpreted within the Bohmian (hydrodynamic) framework [14]. The present formulation provides a complementary theoretical structure in which such flow behaviour emerges from branch-wise organisation induced by gauge coupling.
The results establish branch regularisation as a fundamental mechanism for resolving nonlocal amplitude structures in stationary Bohmian systems, particularly in the presence of electromagnetic gauge fields.
Declaration of Interests
The research is an independent work of the author, self-funded, and declares no conflicts of interest with the employer, IBM Research, Albany, NY, USA. While care has been taken to acknowledge relevant prior work, any inadvertent omissions are unintentional. The author welcomes appropriate references and perspectives to be incorporated in future revisions.
Use of Generative AI
Generative AI tools were used in the preparation of this manuscript solely for language editing, grammar refinement, and equation cross-verification. These tools were not used to generate the scientific results, derivations, or core interpretations presented here. All mathematical checks, technical conclusions, and final editorial decisions remain the sole responsibility of the author.
Appendix Appendix A Bohm–Madelung reduction in cylindrical coordinates
We consider a particle of mass $m$ and charge $e$ in a uniform magnetic field $\bm{B}=B\hat{\bm{z}}$, with vanishing scalar potential. The stationary Schrödinger equation with minimal coupling reads  [1]
$\frac{1}{2m}(\bm{p}-e\bm{A})^{2}\psi=E\psi,$ (A.1)
with the polar decomposition
$\psi=\sqrt{\rho}\,e^{iS/\hbar}.$ (A.2)
Appendix A.1 Coordinate System and Notation
We work in cylindrical coordinates $(r,\theta,z)$ with metric factors
$ds^{2}=dr^{2}+r^{2}d\theta^{2}+dz^{2},\qquad h_{r}=1,\quad h_{\theta}=r,\quad h_{z}=1.$ (A.3)
The vector potential for a uniform magnetic field $B\hat{z}$ is taken in symmetric gauge as
$\mathbf{A}=A_{\theta}\,\hat{\theta}=\frac{Br}{2}\,\hat{\theta}$. (A.4)
We adopt the consistent Bohmian definition of momentum components as the physical gradient components of the action:
$p_{r}=\partial_{r}S,\qquad p_{\theta}=\frac{1}{r}\partial_{\theta}S,\qquad p_{z}=\partial_{z}S.$ (A.5)
The stationary de Broglie–Bohm continuity equation is
$\nabla\cdot\mathbf{J}=0,\qquad\mathbf{J}=\frac{\rho}{m}\,(\nabla S-e\mathbf{A}).$ (A.6)
Appendix A.2 Bohmian Hamiltonian with Minimal Coupling
The stationary de Broglie–Bohm Hamiltonian is
$H_{B}=\frac{1}{2m}(\nabla S-e\mathbf{A})^{2}+Q,$ (A.7)
where the quantum potential is
$Q=-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}.$ (A.8)
Using the cylindrical form of $\nabla S$, we obtain
$\nabla S-e\mathbf{A}=\hat{r}\,p_{r}+\hat{\theta}\left(p_{\theta}-\frac{eBr}{2}\right)+\hat{z}\,p_{z}.$ (A.9)
Hence the Bohmian Hamiltonian becomes
$H_{B}=\frac{1}{2m}\left[p_{r}^{2}+\left(p_{\theta}-\frac{eBr}{2}\right)^{2}+p_{z}^{2}\right]+Q.$ (A.10)
The stationary Bohm equation $H_{B}=E$ with expanded form of $Q$ becomes
$\frac{p_{r}^{2}}{2m}+\frac{p_{\theta}^{2}}{2m}-\frac{eBr}{2m}p_{\theta}+\frac{e^{2}B^{2}r^{2}}{8m}+\frac{p_{z}^{2}}{2m}-\frac{\hbar^{2}}{2m}\left(\frac{R^{\prime\prime}}{R}+\frac{1}{r}\frac{R^{\prime}}{R}+\frac{1}{r^{2}}\frac{\Theta^{\prime\prime}}{\Theta}+\frac{Z^{\prime\prime}}{Z}\right)=E.$ (A.11)
The divergence condition Eq. (A.6) gives us the identity
$\frac{1}{r}\frac{\partial}{\partial r}(r\rho p_{r})+\frac{1}{r}\frac{\partial}{\partial\theta}\left[\rho\left(p_{\theta}-\frac{eBr}{2}\right)\right]+\frac{\partial}{\partial z}(\rho p_{z})=0.$ (A.12)